### v2

# Clutter-Aware Target Detection for ISAC in a Millimeter-Wave Cell-Free Massive MIMO System

Abstract
In this paper, we investigate the performance of an integrated sensing and communication (ISAC) system within a cell-free massive multiple-input multiple-output (MIMO) system. Each access point (AP) operates in the millimeter-wave (mmWave) frequency band. The APs jointly serve the user equipments (UEs) in the downlink while simultaneously detecting a target through dedicated sensing beams, which are directed toward a reconfigurable intelligent surface (RIS). Although the AP-RIS, RIS-target, and AP-target channels have both line-of-sight (LoS) and non-line-of-sight (NLoS) parts, it is assumed only knowledge of the LoS paths is available. A key contribution of this study is the consideration of clutter, which degrades the target detection if not handled. We propose an algorithm to alternatively optimize the transmit power allocation and the RIS phase-shift matrix, maximizing the target signal-to-clutter-plus-noise ratio (SCNR) while ensuring a minimum signal-to-interference-plus-noise ratio (SINR) for the UEs. Numerical results demonstrate that exploiting clutter subspace significantly enhances detection probability, particularly at high clutter-to-noise ratios, and reveal that an increased number of transmit side clusters impairs detection performance. Finally, we highlight the performance gains achieved using a dedicated sensing stream.
Index Terms:
RIS, ISAC, cell-free massive MIMO, mmWave.
I Introduction
Cell-free massive multiple-input multiple-output (MIMO) is a user-centric network infrastructure where a set of distributed access points serve multiple user equipments on the same time-frequency resources using joint processing techniques [1]. Over nearly a decade of research, it has been shown that cell-free massive MIMO improves both spectral and energy efficiency by providing macro diversity and enhanced interference management compared to traditional cellular setups. As a result, it has become one of the key technologies for sixth-generation (6G) wireless networks [2].
Recently, cell-free massive MIMO architecture has also demonstrated potential benefits in integrated sensing and communications (ISAC) [3, 4]. ISAC allows for the efficient use of hardware and spectral resources through integration and coordination gains, compared to separate communication and sensing architectures [5]. Cell-free massive MIMO facilitates multi-static sensing through a central processing unit (CPU), to which the APs are connected via fronthaul links. Beyond cell-free massive MIMO and ISAC, another key 6G technology is reconfigurable intelligent surface (RIS)-aided communications. RISs have attracted significant attention from both academia and industry due to their ability to control challenging radio propagation environments intelligently. By using low-cost passive reflecting elements that induce controllable phase shifts in incoming waveforms, RISs can steer reflected signals toward desired locations while reducing interference at undesired points [6, 7]. Furthermore, RISs have emerged as a promising technology in ISAC applications, enhancing the available spatial degrees-of-freedom. In [8], the sum rate of communication UEs is maximized under a worst-case sensing signal-to-noise ratio (SNR) constraint. An often overlooked aspect in ISAC research is the effect of clutter on target detection, which can severely degrade system performance if not properly addressed. Recent studies, such as [9], illustrate the impact of clutter in monostatic ISAC settings, underscoring the significance of clutter awareness at the receiving APs for reliable target detection.
Refer to captionRISCloud unitAP $r\in\mathcal{R}$TargetClutterAP $t\in\mathcal{T}$UE $k\in\mathcal{K}$$\mathbf{g}_{k,t}$$\mathbf{H}_{t}$$\mathbf{c}$$\mathbf{b}_{r}$ $\overline{\mathbf{H}}_{t}$$\overline{\mathbf{c}}$$\overline{\mathbf{b}}_{r}$ cluster
Figure 1: Depiction of a cell-free massive MIMO ISAC network performing multi-static sensing.
In this paper, we investigate the extent to which an RIS can enhance sensing capabilities when the direct channels between the transmitting APs and the target are obstructed. We also explore the need for precise channel state information (CSI) for target detection within a millimeter-wave (mmWave) channel architecture, as well as the requirement for sensing-specific streams. Our setup, illustrated in Fig. 1, evaluates the system’s ability to detect a potential target at a known location while meeting a minimum signal-to-interference-plus-noise ratio (SINR) for each UE. To this extent, we maximize the target’s signal-to-clutter-and-noise ratio (SCNR) using an alternating optimization (AO) algorithm that jointly optimizes transmit power allocation and the RIS phase-shift matrix. The APs assume a line-of-sight (LoS) structure for the sensing links, justified by the high path loss in mmWave frequencies, which typically makes LoS paths dominant over non-line-of-sight (NLoS) paths. A centralized precoding strategy is employed with regularized zero forcing (RZF) for communication streams. The sensing stream is projected onto the null space of the communication channels, which are assumed to be known at the APs. Assuming knowledge of the clutter subspace, target detection is conducted using a generalized likelihood ratio test (GLRT). The numerical simulations reveal that clutter awareness significantly enhances the detection probability, particularly at high clutter-to-noise ratio (CNR). Furthermore, when sensing links contain many clusters concentrated in a narrow angular range, the LoS assumption becomes less effective, reducing detection probability. The results also highlight the importance of allocating dedicated sensing symbols to improve detection performance.
Notation: $\odot$ represents the Hadamard (element-wise) product. Boldface lowercase and uppercase letters denote vectors and matrices. The trace of the matrix $\mathbf{X}$ is denoted by $\text{Tr}(\mathbf{X})$. $\text{diag}(\mathbf{x})$ represents the stacking of $\mathbf{x}$ on the main diagonal of a matrix. $\mathbf{I}_{N}$ represents an $N\times N$ identity matrix. The notation $\mathcal{CN}(0,\sigma^{2})$ represents circularly symmetric complex Gaussian distribution with variance $\sigma^{2}$. By $\mathcal{L}\setminus\{l\}$, we denote the set obtained by removing the element $l$ from the set $\mathcal{L}$. Given a matrix $\mathbf{M}$, $\mathbf{M}^{\dagger}$ represents its Moore-Penrose pseudoinverse.
II System Model
We consider the RIS-assisted cell-free massive MIMO ISAC network shown in Fig. 1. Here a set $\mathcal{T}$ made of $T$ APs, each equipped with $M$ transmit antennas, serves $K$ single-antenna UEs in the downlink while sensing the potential presence of a target at a known position. The sensing is assumed to be carried out over $\tau$ symbols belonging to the same channel coherence block. A second set of $R$ APs, denoted by $\mathcal{R}$, collects the echos scattered from the target and performs target detection. The choice of geographically separated transmitting and receiving APs is motivated by the implementation challenges of full duplex operation, especially in terms of self interference [10], and the interest in this setup for cell-free massive MIMO communications. We assume that the direct channel between the APs in $\mathcal{T}$ and the target is blocked and, thus, the formers reach the target through a RIS equipped with $N$ reflective elements whose phase shifts are described by the vector $\boldsymbol{\theta}=[\theta_{1},\dots,\theta_{N}]^{\top}$, where $|\theta_{n}|=1$ for $n=1,\dots,N$. Each AP $t\in\mathcal{T}$ is connected to the RIS via the channel $\mathbf{H}_{t}\in\mathbb{C}^{N\times M}$, the reflected path connecting the RIS to the target is denoted by $\mathbf{c}\in\mathbb{C}^{N}$ and the UEs are connected to the APs in $\mathcal{T}$ via the channel $\{\mathbf{g}_{k,t}\}_{\forall k,t}\in\mathbb{C}^{M}$. The receiving APs in $\mathcal{R}$ are connected to the target via the channels $\{\mathbf{b}_{r}\}_{\forall r\in\mathcal{R}}\in\mathbb{C}^{M}$.
II-A Channel models
Each of the previously mentioned channels is modeled in accordance to [11], which accurately captures the spatial sparsity that characterizes mmWave channels. Each channel consists of contributions from $C_{i}$ scattering clusters, i.e.,
$\mathbf{b}_{r}=\underbrace{\sqrt{\beta^{2}_{0,r}}\hskip 2.84526pt\mathbf{a}_{M}(\boldsymbol{\omega}_{0,r}^{\textrm{RX}})}_{\overline{\mathbf{b}}_{r}}+\sqrt{\frac{1}{C_{1}}}\sum_{n=1}^{C_{1}}\alpha_{n,r}\hskip 1.42262pt\mathbf{a}_{M}(\boldsymbol{\omega}_{n,r}^{\textrm{RX}}),$ (1)
$\mathbf{c}=\underbrace{\sqrt{\beta_{0}^{2}}\hskip 2.84526pt\mathbf{a}_{N}(\boldsymbol{\omega}_{0}^{\textrm{Tg}})}_{\overline{\mathbf{c}}}+\sqrt{\frac{1}{C_{2}}}\sum_{n=1}^{C_{2}}\alpha_{n}\hskip 1.42262pt\mathbf{a}_{N}(\boldsymbol{\omega}_{n}^{\textrm{Tg}}),$ (2)
$\mathbf{g}_{k,t}=\sqrt{\frac{1}{C_{3}}}\sum_{n=1}^{C_{3}}\alpha_{n,k,t}\hskip 1.42262pt\mathbf{a}_{M}(\boldsymbol{\omega}_{n,k,t}^{\textrm{UE}}),$ (3)
$\mathbf{H}_{t}=\overbrace{\sqrt{\beta_{0,t}^{2}}\hskip 2.84526pt\mathbf{a}_{N}(\boldsymbol{\omega}_{0,t}^{\textrm{RIS}})\mathbf{a}_{M}^{\top}(\boldsymbol{\omega}_{0,t}^{\textrm{TX}})}^{\overline{\mathbf{H}}_{t}}$
[…]
where the overline superscript indicates the LoS paths and the other terms correspond to the NLoS parts. The $\beta_{0,r}^{2}$, $\beta_{0}^{2}$, and $\beta_{0,t}^{2}$ are the corresponding LoS channel gains. We now remove the additional subscripts and superscripts except for $n$ as the following consideration applies regardless of the additional subscripts. Here, $\alpha_{n}$ is the complex channel gain of the $n$-th cluster. We assume each cluster scatters a sufficient number of rays such that $\alpha_{n}\sim\mathcal{CN}(0,\beta_{n}^{2})$ and the fading is mutually independent. The coefficients $\beta_{n}^{2}$ represent the path loss associated with the respective clusters. The geometry of the $n$-th cluster is described by $\boldsymbol{\omega}_{n}=[\psi_{n},\phi_{n}]$, where $\psi_{n}$ is the azimuth angle and $\phi_{n}$ is the elevation angle. Each AP is equipped with a uniform planar square array with a half-wavelength vertical and horizontal inter-element spacing such that the entries of $M$-dimensional steering vector are defined as
$[\mathbf{a}_{M}(\boldsymbol{\omega}_{n})]_{m_{h}m_{v}}=e^{-j\pi(m_{h}\sin(\psi_{n})\cos(\phi_{n})+m_{v}\sin(\phi_{n}))}$ (5)
for $m_{h},m_{v}=0,\dots,\sqrt{M}-1$. We assume that perfect CSI for $\{\mathbf{g}_{k,t}\}_{\forall k,t}$ is available at the APs, as this is easily obtainable through well-known estimation techniques. As for the sensing channels, the transmit APs only know the position of the RIS and the target and, thus, the corresponding LoS channels.
II-B Communication observation model
We assume that the APs are equipped with fully digital beamforming hardware: $\mathbf{f}_{l,t}$ is the precoding vector for stream $l$ and AP $t$, and $\mathbf{F}_{t}=[\mathbf{f}_{1,t},\ldots,\mathbf{f}_{L,t}]\in\mathbb{C}^{M\times L}$, where $L=K+1$. Here, we allocate one beam to each UE and an additional one, denoted by the index $L$, for sensing purposes, bringing the total number of digital beams to $L=K+1$. This choice is motivated by the fact that the UEs and target typically are at very different positions, and catering solely to the communication UEs might drastically reduce the system’s sensing performance.
The received downlink signal at UE $k$ during timeslot $\iota$ is defined as
$y_{k}[\iota]$ $=\sum_{t\in\mathcal{T}}\mathbf{g}_{k,t}^{\mathsf{H}}\mathbf{s}_{t}[\iota]+w_{k}[\iota]=\sum_{t\in\mathcal{T}}\Big{(}\underbrace{\sqrt{\rho_{k}}\mathbf{g}_{k,t}^{\mathsf{H}}\mathbf{f}_{k,t}x_{k}[\iota]}_{\text{desired signal}}$ (6)
$+\underbrace{\sum_{k^{\prime}\in\mathcal{K}\setminus\{k\}}\sqrt{\rho_{k^{\prime}}}\mathbf{g}_{k,t}^{\mathsf{H}}\mathbf{f}_{k^{\prime},t}x_{k^{\prime}}[\iota]}_{\text{multi-user interf. (MUI)}}+\underbrace{\sqrt{\rho_{L}}\mathbf{g}_{k,t}^{\mathsf{H}}\mathbf{f}_{L,t}x_{L}[\iota]}_{\text{sensing interf.}}\Big{)}+w_{k}[\iota],$
with $\mathbf{s}_{t}[\iota]=\mathbf{F}_{t}\mathbf{X}[\iota]\boldsymbol{\rho}$, where the matrix $\mathbf{X}[\iota]=\text{diag}(x_{1}[\iota],\ldots,x_{L}[\iota])$ is a diagonal matrix containing the downlink communication symbols and sensing symbol transmitted during timeslot $\iota$, where $\mathbb{E}\{|x_{k}[\iota]|^{2}\}=1$. The vector $\boldsymbol{\rho}=[\sqrt{\rho_{1}},\dots,\sqrt{\rho_{L}}]^{\top}$ contains the square root of the power allocated to each beam. Finally, $w_{k}[\iota]\sim\mathcal{CN}(0,\sigma_{k}^{2})$. In line with the centralized operation of cell-free massive MIMO, the communication precoding vectors are based on the transmit APs’ CSI [3] while the sensing precoder is based on the assumed LoS structure of the sensing channels and the known target location. To this end, we define $\mathbf{g}_{k}=[\mathbf{g}_{k,1}^{\top},\dots,\mathbf{g}_{k,T}^{\top}]^{\top}\in\mathbb{C}^{MT}$ and $\mathbf{f}_{k}=[\mathbf{f}_{k,1}^{\top},\dots,\mathbf{f}_{k,T}^{\top}]^{\top}\in\mathbb{C}^{MT}$. The communication precoding vectors for each UE are chosen using RZF as
$\mathbf{f}_{k}=\frac{\left(\sum_{i\in\mathcal{K}}\mathbf{g}_{i}\mathbf{g}_{i}^{\mathsf{H}}+\lambda\mathbf{I}_{MT}\right)^{-1}\mathbf{g}_{k}}{\left\lVert\left(\sum_{i\in\mathcal{K}}\mathbf{g}_{i}\mathbf{g}_{i}^{\mathsf{H}}+\lambda\mathbf{I}_{MT}\right)^{-1}\mathbf{g}_{k}\right\rVert},\leavevmode\nobreak\ k=1,\dots,K,$ (7)
where $\lambda$ is a regularization parameter and this parameter allows the system to implement a trade-off between interference and noise suppression. For the sensing stream, let us define the matrix $\mathbf{G}=[\mathbf{g}_{1},\dots,\mathbf{g}_{K}]\in\mathbb{C}^{MT\times K}$ with all the UE channels. Then the sensing stream precoder is defined as
$\mathbf{f}_{L,t}=\frac{\left(\mathbf{I}_{MT}-\mathbf{G}\left(\mathbf{G}^{\mathsf{H}}\mathbf{G}\right)^{\dagger}\mathbf{G}^{\mathsf{H}}\right)\overline{\mathbf{h}}}{\left\lVert\left(\mathbf{I}_{MT}-\mathbf{G}\left(\mathbf{G}^{\mathsf{H}}\mathbf{G}\right)^{\dagger}\mathbf{G}^{\mathsf{H}}\right)\overline{\mathbf{h}}\right\rVert},$ (8)
where $\overline{\mathbf{h}}=[\overline{\mathbf{h}}_{1}^{\top},\dots,\overline{\mathbf{h}}_{T}^{\top}]^{\top}$ and $\overline{\mathbf{h}}_{t}=\overline{\mathbf{H}}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\overline{\mathbf{c}}$ with $\boldsymbol{\Theta}=\text{diag}(\boldsymbol{\theta})$. We have projected $\overline{\mathbf{h}}$ onto the null space of the communication UEs’ channels to null the interference caused to the UEs. The communication performance is represented by the UEs’ spectral efficiency (SE), defined as ${\rm{SE}}_{k}=\log_{2}(1+{\rm{SINR}}_{k})$: under the assumption of perfect CSI on the UEs channels, the SINR is defined as
${\rm{SINR}}_{k}=\frac{|{\mathbf{g}}_{k}^{\mathsf{H}}\mathbf{f}_{k}\sqrt{\rho_{k}}|^{2}}{\sum_{k^{\prime}\in\mathcal{K}/k}|{\mathbf{g}}_{k}^{\mathsf{H}}\mathbf{f}_{k^{\prime}}\sqrt{\rho_{k^{\prime}}}|^{2}+|{\mathbf{g}}_{k}^{\mathsf{H}}\mathbf{f}_{L}\sqrt{\rho_{L}}|^{2}+\sigma_{k}^{2}}.$ (9)
II-C Sensing observation model
The two-way sensing channel between transmit AP $t$ and receive AP $r$ is defined as $\mathbf{E}_{t,r}=\mathbf{b}_{r}\mathbf{h}_{t}^{\mathsf{H}}\triangleq\mathbf{b}_{r}\left(\mathbf{H}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\mathbf{c}\right)^{\mathsf{H}}.$ The sensing observation of AP $r$ at timeslot $\iota$ is
$\mathbf{y}_{r}[\iota]=\sum_{t\in\mathcal{T}}\xi_{t,r}\mathbf{E}_{t,r}\mathbf{s}_{t}[\iota]+\mathbf{z}_{r}[\iota]+\mathbf{n}_{r}[\iota],$ (10)
where $\xi_{t,r}$ is the radar cross section (RCS) associated with the sensing path between transmit AP $t$ and receive AP $r$. We adopt the Swerling-I model, meaning that the RCS assumes only one value during the transmission and $\xi_{t,r}\sim\mathcal{CN}(0,\delta_{t,r}^{2})$. We assume that $\{\delta_{t,r}^{2}\}_{\forall t,r}$ are known and that RCSs belonging to different transmit-receiving pairs are statistically uncorrelated, that is
$\mathbb{E}\left\{\xi_{a,r}\xi_{a^{\prime},r^{\prime}}^{*}\right\}=\begin{cases}\delta_{a,r}^{2},&{\rm{if}}\leavevmode\nobreak\ a=a^{\prime},\leavevmode\nobreak\ r=r^{\prime},\\ 0,&\rm{otherwise}.\end{cases}$ (11)
In (10), $\mathbf{n}_{r}[\iota]$ is the receiver noise of AP $r$, assumed to have independent $\mathcal{CN}(0,\sigma^{2})$ entries. The vector $\mathbf{z}_{r}[\iota]$ represents the clutter as observed from AP $r$. The clutter vector is modeled as $\mathbf{z}_{r}[\iota]\sim\mathcal{CN}(\boldsymbol{0},\delta_{\mathbf{z}}^{2}\mathbf{R}_{r})$, where $\delta_{\mathbf{z}}^{2}$ is the power of the clutter cross section, which also includes the path loss effects, thus defining the CNR as $\textrm{CNR}=\delta_{\mathbf{z}}^{2}/\sigma^{2}$. The normalized clutter covariance matrix with trace being equal to $M$ is denoted by $\mathbf{R}_{r}$. This matrix is unknown, but its eigenspace is assumed to be known later in the detector design. For the sake of mathematical tractability, we assume that the clutter observations belonging to different APs and/or different timeslots are statistically uncorrelated. Under the LoS assumption for the channels involved in sensing, the target’s SCNR is defined as
$\overline{\textrm{SCNR}}=\frac{\mathbb{\sum}_{\iota=1}^{\tau}\mathbb{\sum}_{r\in\mathcal{R}}\sum_{a\in\mathcal{T}}\delta_{a,r}^{2}\left\lVert\overline{\mathbf{E}}_{a,r}\mathbf{s}_{a}[\iota]\right\rVert^{2}}{R\tau M\left(\sigma^{2}+\delta_{\mathbf{z}}^{2}\right)},$ (12)
where $\overline{\mathbf{E}}_{a,r}=\overline{\mathbf{b}}_{r}\left(\overline{\mathbf{H}}_{a}^{\mathsf{H}}\boldsymbol{\Theta}\overline{\mathbf{c}}\right)^{\mathsf{H}}$
III Sensing performance optimization
The system aims at jointly maximizing the sensing SCNR (12) while guaranteeing a minimum SINR to all the communication UEs. The optimization problem is stated as follows:
$\underset{\boldsymbol{\theta},\boldsymbol{\rho}}{\mathrm{maximize}}\,\,$ $\leavevmode\nobreak\ \overline{\textrm{SCNR}}$ (13a)
subject to $\textrm{SINR}_{k}\geq\gamma_{k},\quad\forall k\in\mathcal{K},$ (13b)
$\sum_{l=1}^{L}\left\lVert\mathbf{f}_{l,t}\right\rVert^{2}\rho_{l}\leq P_{t},\quad\forall t\in\mathcal{T},$ (13c)
$|\theta_{n}|=1,\quad n=1,\dots,N,$ (13d)
where $\gamma_{k}$ is the SINR threshold of UE $k$ and $P_{t}$ is the available transmit power of the $t$-th AP. This problem is non-convex, mainly due to the coupling between the optimization variables and the non-convex unitary modulus constraints. In this section, we will devise an AO algorithm, which switches between optimizing the power allocation policy and the RIS phase shifts until convergence is achieved.
III-A Power allocation optimization
In this subsection, we introduce the first subproblem of the AO algorithm, where we optimize the power allocation for a fixed RIS configuration. We first reformulate (12) as
$\overline{\textrm{SCNR}}=\boldsymbol{\rho}^{\top}\Re(\mathbf{A})\boldsymbol{\rho},$ (14)
where
$\mathbf{A}=\frac{\sum_{\iota=1}^{\tau}\mathbb{\sum}_{r\in\mathcal{R}}\sum_{a\in\mathcal{T}}\delta_{a,r}^{2}\mathbf{X}^{\mathsf{H}}[\iota]\mathbf{F}_{a}^{\mathsf{H}}\overline{\mathbf{E}}_{a,r}^{\mathsf{H}}\overline{\mathbf{E}}_{a,r}\mathbf{F}_{a}\mathbf{X}[\iota]}{R\tau M\left(\sigma^{2}+\delta_{\mathbf{z}}^{2}\right)}.$ (15)
The set of constraints (13b) can be reformulated as second order cone (SOC) constraints in terms of $\boldsymbol{\rho}$ as
$\left\lVert\begin{bmatrix}|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{1}|\sqrt{\rho_{1}}&\dots&\underbrace{0}_{k}&\dots&|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{L}|\sqrt{\rho_{L}}&\sigma_{k}\end{bmatrix}\right\rVert$
$\leq\frac{|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{k}|\sqrt{\rho_{k}}}{\sqrt{\gamma_{k}}}.$ (16)
The power constraint in (13c) can also be rewritten in SOC form as
$\left\lVert\text{diag}\left(\left\lVert\mathbf{f}_{1,t}\right\rVert,\dots,\left\lVert\mathbf{f}_{L,t}\right\rVert\right)\boldsymbol{\rho}\right\rVert\leq\sqrt{P_{t}}.$ (17)
The first subproblem of the AO SCNR-maximizing algorithm is defined as
$\underset{\boldsymbol{\rho}}{\mathrm{maximize}}\,\,$ $\leavevmode\nobreak\ \boldsymbol{\rho}^{\top}\Re(\mathbf{A})\boldsymbol{\rho}$ (18a)
[…]
This problem can be solved using the convex-concave procedure outlined in [3].
III-B RIS phase-shift optimization
We now move on to the second part of the AO, where we optimize $\boldsymbol{\theta}$ when the power variables are fixed. The SCNR can be reformulated as
$\overline{\textrm{SCNR}}=\boldsymbol{\theta}^{\mathsf{H}}\mathbf{Q}\boldsymbol{\theta},$ (19)
where
$\mathbf{Q}=$ $\frac{1}{R\tau M\left(\sigma^{2}+\delta_{\mathbf{z}}^{2}\right)}\sum_{\iota=1}^{\tau}\sum_{r\in\mathcal{R}}\sum_{a\in\mathcal{T}}\Big{(}\delta_{a,r}^{2}\overline{\mathbf{c}}\overline{\mathbf{b}}_{r}^{\mathsf{H}}\overline{\mathbf{b}}_{r}\overline{\mathbf{c}}^{\mathsf{H}}\Big{)}^{\top}$
$\quad\quad\odot\left(\overline{\mathbf{H}}_{a}\mathbf{s}_{a}[\iota]\mathbf{s}_{a}^{\mathsf{H}}[\iota]\overline{\mathbf{H}}_{a}^{\mathsf{H}}\right).$ (20)
We have used $\text{Tr}(\mathbf{A}\mathbf{X}\mathbf{B}\mathbf{X}^{\mathsf{H}})=\text{vec}^{\mathsf{H}}(\mathbf{X})\left(\mathbf{B}^{\top}\otimes\mathbf{A}\right)\text{vec}(\mathbf{X})$ and the following lemma.
Lemma 1.
Let $\mathbf{x}\in\mathbb{C}^{N}$ and $\mathbf{X}=\text{diag}(\mathbf{x})$. For any $\mathbf{A}\in\mathbb{C}^{N\times N}$ and $\mathbf{B}\in\mathbb{C}^{N\times N}$, it holds that
$\text{vec}^{\mathsf{H}}(\mathbf{X})(\mathbf{A}\otimes\mathbf{B})\text{vec}(\mathbf{X})=\mathbf{x}^{\mathsf{H}}(\mathbf{A}\odot\mathbf{B})\mathbf{x}.$ (21)
Proof: Let $\text{vec}(\mathbf{X})=\sum_{n=1}^{N}x_{n}\mathbf{e}_{n}\otimes\mathbf{e}_{n}$, where $\mathbf{e}_{n}$ is the $n$-th column of $\mathbf{I}_{N}$. Then, the right-hand side of (21) can be written as $\left(\sum_{n=1}^{N}x_{n}\mathbf{e}_{n}\otimes\mathbf{e}_{n}\right)^{\mathsf{H}}(\mathbf{A}\otimes\mathbf{B})\left(\sum_{m=1}^{N}x_{m}\mathbf{e}_{m}\otimes\mathbf{e}_{m}\right)\overset{(a)}{=}\sum_{n=1}^{N}\sum_{m=1}^{N}x_{n}x_{m}^{\mathsf{H}}(\mathbf{e}_{n}^{\mathsf{H}}\mathbf{A}\mathbf{e}_{m})(\mathbf{e}_{n}^{\mathsf{H}}\mathbf{B}\mathbf{e}_{m})=\sum_{n=1}^{N}\sum_{m=1}^{N}x_{n}x_{m}^{\mathsf{H}}[\mathbf{A}]_{n,m}[\mathbf{B}]_{n,m}=\mathbf{x}^{\mathsf{H}}(\mathbf{A}\odot\mathbf{B})\mathbf{x}$, where in $(a)$ we used mixed-product property of the Kronecker product.
The RIS phase-shift optimization problem under consideration is still non-convex. We are trying to maximize a quadratic form of $\boldsymbol{\theta}$. We then employ the minorization-maximization (MM) algorithm to maximize a convex lower bound on (19) [12]. More specifically, a convex local lower bound can be found by applying the successive convex approximation [13] around the local point $\boldsymbol{\theta}^{(s)}$ at iteration $s$. The objective function can be lower bounded by
$\boldsymbol{\theta}^{\mathsf{H}}\mathbf{Q}\boldsymbol{\theta}\geq 2\Re\left(\boldsymbol{\theta}^{(s)\mathsf{H}}\mathbf{Q}^{(+)}\boldsymbol{\theta}\right)+\boldsymbol{\theta}^{\mathsf{H}}\mathbf{Q}^{(-)}\boldsymbol{\theta}-\boldsymbol{\theta}^{(s)\mathsf{H}}\mathbf{Q}^{(+)}\boldsymbol{\theta}^{(s)},$ (22)
where $\mathbf{Q}^{(+)}$ and $\mathbf{Q}^{(-)}$ represent the positive and negative semidefinite parts of $\mathbf{Q}$. We can now define the second subproblem of the AO optimization as
$\underset{\boldsymbol{\theta},\alpha\geq 0}{\textrm{maximize}}$ $\leavevmode\nobreak\ \alpha$ (23a)
subject to $\alpha\leq 2\Re\left(\boldsymbol{\theta}^{(s)\mathsf{H}}\mathbf{Q}^{(+)}\boldsymbol{\theta}\right)+\boldsymbol{\theta}^{\mathsf{H}}\mathbf{Q}^{(-)}\boldsymbol{\theta}$ (23b)
$\quad\quad-\boldsymbol{\theta}^{(s)\mathsf{H}}\mathbf{Q}^{(+)}\boldsymbol{\theta}^{(s)},$
$|\theta_{n}|\leq 1,\quad n=1,\dots,N,$ (23c)
where $\alpha$ is an auxiliary variable. Note that we have relaxed the unit modulus constraint, as it is observed that the solution satisfies this constraint with equality anyway. This is a convex problem and can be solved at each iteration until convergence is achieved. The steps of the AO algorithm are outlined in Algorithm 1.
1:Initialize: Generate $\boldsymbol{\theta}^{(0)}$ randomly, $v\leftarrow 0$
2:repeat
3:   Compute the precoding codebooks with $\boldsymbol{\theta}=\boldsymbol{\theta}^{(v)}$
4:   Obtain $\boldsymbol{\rho}^{(v+1)}$ by solving (18)
5:   Obtain $\boldsymbol{\theta}^{(v+1)}$ by solving (23) iteratively
6:   $v\leftarrow v+1$
7:until Convergence
8:Output: $\boldsymbol{\theta}^{\text{opt}},\boldsymbol{\rho}^{\text{opt}}$
Algorithm 1 SCNR maximizing AO algorithm
IV GLRT Detector
One of the main metrics used to assess the sensing performance is the probability of correct detection, implemented through a GLRT by extending the detector in [9] to a cell-free massive MIMO system based on the optimization procedure developed in the previous section. We operate under the assumption that the transmit signals are known by the receiving APs, thanks to the cell-free architecture. We define the vector containing the combined sensing observation of all receiving APs during timeslot $\iota$ as
$\mathbf{y}[\iota]=[\mathbf{y}_{1}^{\top}[\iota],\dots,\mathbf{y}_{R}^{\top}[\iota]]^{\top}=\mathbf{V}[\iota]\boldsymbol{\xi}+\mathbf{U}\mathbf{w}[\iota]+\mathbf{n}[\iota],$ (24)
$\mathbf{V}[\iota]=\textrm{blkdiag}(\mathbf{V}_{1}[\iota],\dots,\mathbf{V}_{R}[\iota])\in\mathbb{C}^{MR\times TR},$ (25)
$\mathbf{V}_{r}[\iota]=\left[\mathbf{E}_{1,r}\mathbf{s}_{1}[\iota],\dots,\mathbf{E}_{T,r}\mathbf{s}_{T}[\iota]\right]\in\mathbb{C}^{M\times T},$ (26)
$\mathbf{n}[\iota]=[\mathbf{n}_{1}^{\top}[\iota],\dots,\mathbf{n}_{R}^{\top}[\iota]]^{\top},\leavevmode\nobreak\ \boldsymbol{\xi}=[\xi_{1,1},\dots,\xi_{T,R}]^{\top},$ (27)
$\mathbf{U}=\textrm{blkdiag}\left(\mathbf{U}_{1},\dots,\mathbf{U}_{R}\right),\mathbf{w}[\iota]=[\mathbf{w}_{1}^{\top}[\iota],\dots,\mathbf{w}_{R}^{\top}[\iota]]^{\top}.$ (28)
The clutter observed by AP $r$ is assumed to exist in a low-rank subspace [14] spanned by the columns of the semi-unitary matrix $\mathbf{U}_{r}\in\mathbb{C}^{M\times i}$, with $i$ being the subspace dimension, obtained from the eigenspace of $\delta_{\mathbf{z}}^{2}\mathbf{R}_{r}$. AP $r$’s clutter is then expressed as $\mathbf{U}_{r}\mathbf{w}_{r}[\iota]$, where the distribution of $\mathbf{w}_{r}[\iota]$ is unknown. The previous observations are generated using the true channels; however, the system does not have perfect CSI on the sensing channels and thus assumes a LoS structure. Therefore, the binary hypotheses being tested are
$\mathcal{H}_{0}:\mathbf{y}[\iota]=\overline{\mathbf{V}}[\iota]\boldsymbol{\xi}+\mathbf{U}\mathbf{w}[\iota]+\mathbf{n}[\iota],\leavevmode\nobreak\ \iota=1,\dots,\tau,$ (29a)
$\mathcal{H}_{1}:\mathbf{y}[\iota]=\mathbf{U}\mathbf{w}[\iota]+\mathbf{n}[\iota],\leavevmode\nobreak\ \iota=1,\dots,\tau,$ (29b)
where $\overline{\mathbf{V}}$ is computed as $\mathbf{V}$ by using $\overline{\mathbf{E}}_{t,r}$. The adoption of a GLRT method is motivated by the unknown distribution of $\mathbf{w}[\iota]$, however we know that $\mathbf{y}_{\mathcal{H}_{0}}[\iota]-\mathbf{U}\mathbf{w}[\iota]\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{MR})$ and $\mathbf{y}_{\mathcal{H}_{1}}[\iota]-\mathbf{U}\mathbf{w}[\iota]\sim\mathcal{CN}(\mathbf{0},\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR})$, where $\mathbf{R}[\iota]=\mathbb{E}\left\{\overline{\mathbf{V}}[\iota]\boldsymbol{\xi}\boldsymbol{\xi}^{\mathsf{H}}\overline{\mathbf{V}}^{\mathsf{H}}[\iota]\right\}=\overline{\mathbf{V}}[\iota]\text{diag}(\delta^{2}_{1,1},\dots,\delta^{2}_{T,R})\overline{\mathbf{V}}^{\mathsf{H}}[\iota]$. The GLRT test is then defined as
$\frac{\prod_{\iota=1}^{\tau}\max_{\mathbf{w}[\iota]}f_{\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}_{1}}\left(\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}_{1}\right)}{\prod_{\iota=1}^{\tau}\max_{\mathbf{w}[\iota]}f_{\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}_{0}}\left(\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}_{0}\right)}\lessgtr^{\mathcal{H}_{0}}_{\mathcal{H}_{1}}\lambda_{d},$ (30)
where $f_{\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}}$ denote the probability density function. Following the derivation in [9], the GLRT test can be rewritten as
$\sum_{\iota=1}^{\tau}\mathbf{y}^{\mathsf{H}}[\iota]\mathbf{T}[\iota]\mathbf{y}[\iota]\lessgtr^{\mathcal{H}_{0}}_{\mathcal{H}_{1}}\underbrace{\ln(\lambda_{d})-\ln\left(\varsigma\right)}_{\widetilde{\lambda}_{d}}$ (31)
$\mathbf{T}[\iota]=\boldsymbol{\Xi}[\iota]\mathbf{U}\Big{(}\mathbf{U}^{\mathsf{H}}\boldsymbol{\Xi}[\iota]\mathbf{U}\Big{)}^{-1}\mathbf{U}^{\mathsf{H}}\boldsymbol{\Xi}[\iota]$
$\quad\quad+\sigma^{-2}(\mathbf{I}_{MR}-\mathbf{U}\mathbf{U}^{\mathsf{H}})-\boldsymbol{\Xi}[\iota]$ (32)
$\boldsymbol{\Xi}[\iota]=(\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR})^{-1}$ (33)
where $\varsigma=\det\left(\sigma^{2}\mathbf{I}_{MR}\right)^{\tau}/{\prod_{\iota=1}^{\tau}\det\left(\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR}\right)}$. $\widetilde{\lambda}_{d}$ is a revised threshold whose value should be chosen according to the desired false alarm probability, and it can be computed through Monte Carlo trials.
5050505010010010010015015015015000101010102020202030303030X (m)Y (m)$\mathcal{T}$$\mathcal{R}$
Figure 2: 2D locations in the simulated scenario.
V Numerical Results
In this section, the numerical simulation results are provided. A two-dimensional view of the simulated scenario is provided in Fig. 2. The locations of the transmit APs in $\mathcal{T}$, the receiving ones in $\mathcal{R}$, and the UEs are randomly generated within the respective bounding boxes, and their $z$ coordinates are equal to $10$ m, $10$ m, and $1$ m, respectively. The RIS and target position are assumed to be fixed and their $z$ coordinates are equal to $15$ m and $3$ m. The power budget for each transmit AP is $2$ dBW, with a bandwidth of $B=1$ MHz, the noise power is equal to $-204+10\log_{10}(B)$ dBW. The channel’s path loss follows the 3GPP urban microcell model defined in [15]. The clutter covariance matrix $\mathbf{R}_{r}$ is modeled according to the local scattering model in [16] with six clusters centered around the $\boldsymbol{\omega}_{0,r}^{\textrm{RX}}\pm[20^{\circ}$, $10^{\circ}]$. We have $T=5$ transmit APs and $R=2$ receiving APs, each equipped with $M=36$ antennas and serving $K=5$ UEs. We want to guarantee an SINR of $\gamma_{k}=3$ dB to all UEs. Unless otherwise specified, the RIS is equipped with $N=64$ elements, $C_{1}=C_{2}=C_{3}=C_{4}=2$, and the angles are uniformly distributed within an angle spread of $\sigma_{\psi_{n}}=10^{\circ}$ in azimuth and $\sigma_{\phi_{n}}=10^{\circ}$ in elevation around the LoS angles between transmitter and receiver, denoted by $\psi_{0}\leavevmode\nobreak\ (\phi_{0})$. The results shown hereafter are averaged over $10$ random realizations of the AP positions and UE positions. For each of these realizations, the probability of detection is computed over $\tau=5$ timeslots and $1000\tau$ noise and clutter realizations. The false alarm probability is set to $10^{-4}$.
[…]
Figure 3: Probability of detection with respect to the RCS power with a CNR of $20$ dB.
[…]
Figure 4: Probability of detection with respect to the RCS power with a CNR of $40$ dB.
V-A Clutter awareness and cluster number’s impact
We first analyze the impact of the clutter awareness of the receiving APs on the detection probability and how much the LoS assumed structure of the channels affects the system’s detection performances when there are many clusters concentrated within a small angular region. Fig. 3 shows this probability of detection vs. the RCS power, where we assumed that $\delta_{1,1}^{2}=\dots=\delta_{T,R}^{2}=\delta^{2}$ for a CNR of $20$ dB. We can see that a clutter-aware system outperforms a clutter-unaware one: The clutter unaware dectector is obtained from (29) by removing the term $\mathbf{U}\mathbf{w}[\iota]$, thus obtaining $\mathbf{T}[\iota]=\sigma^{-2}\mathbf{I}_{MR}-\boldsymbol{\Xi}[\iota]$. On the other hand, when $C_{2}=C_{3}=C_{4}=20$, here denoted by $C_{234}$, the LoS assumption does not hold as much, thus generating a performance decrease. Our optimized model greatly outperforms a benchmark approach, consisting of a random RIS phase-shift matrix and an equal power allocation between the beams. Fig. 4 shows what happens when the CNR rises to $40$ dB. We see that clutter awareness rises in importance, widening the gap between the clutter-aware and unaware detectors.
[…]
Figure 5: Impact of a dedicated sensing stream onto the probability of detection with a CNR of $20$ dB.
V-B Do we need dedicated sensing symbols?
Given the sparse nature of the mmWave channels, one might question whether or not a dedicated sensing stream is necessary and to what extent the system benefits from it. To gain more insight into this, we have analyzed a system without sensing-dedicated precoding. Fig. 5 compares the detection performance of systems with and without dedicated sensing precoding, both in the presence and in the absence of clutter awareness. We can immediately see that removing the sensing streams implies a performance degradation. Interestingly, the system without sensing dedicated precoding seems to suffer less from the lack of clutter awareness at a low RCS value.
VI Conclusions
We have investigated a cell-free massive MIMO ISAC network jointly serving multiple communication UEs while detecting a potential target through an RIS in a multi-static fashion. The system wants to maximize its sensing performance while preserving a minimum SINR for the communication UEs. To this end, we optimize the power allocation of a precoding codebook jointly with the RIS phase-shift matrix and we derive the sufficient statistics for the target detector. Numerical results show that clutter awareness at the receiving APs plays a crucial role in the system’s detection performance and that including a dedicated sensing stream benefits the system. We see that the mmWave channel structure allows the system to assume LoS channels and still achieve good performances when the channel is composed of a high number of clusters concentrated in a small angular range.

### v3

@@
 Each access point (AP) operates in the millimeter-wave (mmWave) frequency band.
-The APs jointly serve the user equipments (UEs) in the downlink while simultaneously detecting a target through dedicated sensing beams, which are directed toward a reconfigurable intelligent surface (RIS).
-Although the AP-RIS, RIS-target, and AP-target channels have both line-of-sight (LoS) and non-line-of-sight (NLoS) parts, it is assumed only knowledge of the LoS paths is available.
-A key contribution of this study is the consideration of clutter, which degrades the target detection if not handled.
+The APs jointly serve the user equipments (UEs) in the downlink while simultaneously detecting a target through dedicated sensing beams directed toward a reconfigurable intelligent surface (RIS).
+Although the AP-RIS, RIS-target, and AP-target channels have both line-of-sight (LoS) and non-line-of-sight (NLoS) parts, only knowledge of the LoS paths is assumed to be available.
+A key contribution of this study is the consideration of clutter, which degrades the target detection performance if not handled.
 We propose an algorithm to alternatively optimize the transmit power allocation and the RIS phase-shift matrix, maximizing the target signal-to-clutter-plus-noise ratio (SCNR) while ensuring a minimum signal-to-interference-plus-noise ratio (SINR) for the UEs.
@@
 Cell-free massive multiple-input multiple-output (MIMO) is a user-centric network infrastructure where a set of distributed access points serve multiple user equipments on the same time-frequency resources using joint processing techniques [1].
-Over nearly a decade of research, it has been shown that cell-free massive MIMO improves both spectral and energy efficiency by providing macro diversity and enhanced interference management compared to traditional cellular setups.
-As a result, it has become one of the key technologies for sixth-generation (6G) wireless networks [2].
+Over nearly a decade of research, it has been shown that cell-free massive MIMO improves both spectral and energy efficiency by providing macro-diversity and enhanced interference management compared to traditional cellular setups.
+As a result, it has become one of the key technology components considered for sixth-generation (6G) wireless networks [2].
 Recently, cell-free massive MIMO architecture has also demonstrated potential benefits in integrated sensing and communications (ISAC) [3, 4].
@@
 Beyond cell-free massive MIMO and ISAC, another key 6G technology is reconfigurable intelligent surface (RIS)-aided communications.
-RISs have attracted significant attention from both academia and industry due to their ability to control challenging radio propagation environments intelligently.
-By using low-cost passive reflecting elements that induce controllable phase shifts in incoming waveforms, RISs can steer reflected signals toward desired locations while reducing interference at undesired points [6, 7].
-Furthermore, RISs have emerged as a promising technology in ISAC applications, enhancing the available spatial degrees-of-freedom.
+RISs have emerged as a promising technology in ISAC applications, enhancing the available spatial degrees of freedom.
 In [6], the sum rate of communication UEs is maximized under a worst-case sensing signal-to-noise ratio (SNR) constraint.
@@
 Figure 1: Depiction of a cell-free massive MIMO ISAC network performing multi-static sensing.
-In this paper, we investigate the extent to which an RIS can enhance sensing capabilities when the direct channels between the transmitting APs and the target are obstructed.
-We also explore the need for precise channel state information (CSI) for target detection within a millimeter-wave (mmWave) channel architecture, as well as the requirement for sensing-specific streams.
-Our setup, illustrated in Fig.
-1, evaluates the system’s ability to detect a potential target at a known location while meeting a minimum signal-to-interference-plus-noise ratio (SINR) for each UE.
-To this extent, we maximize the target’s signal-to-clutter-and-noise ratio (SCNR) using an alternating optimization (AO) algorithm that jointly optimizes transmit power allocation and the RIS phase-shift matrix.
-The APs assume a line-of-sight (LoS) structure for the sensing links, justified by the high path loss in mmWave frequencies, which typically makes LoS paths dominant over non-line-of-sight (NLoS) paths.
-A centralized precoding strategy is employed with regularized zero forcing (RZF) for communication streams.
-The sensing stream is projected onto the null space of the communication channels, which are assumed to be known at the APs.
-Assuming knowledge of the clutter subspace, target detection is conducted using a generalized likelihood ratio test (GLRT).
-The numerical simulations reveal that clutter awareness significantly enhances the detection probability, particularly at high clutter-to-noise ratio (CNR).
-Furthermore, when sensing links contain many clusters concentrated in a narrow angular range, the LoS assumption becomes less effective, reducing detection probability.
-The results also highlight the importance of allocating dedicated sensing symbols to improve detection performance.
-Notation: $\odot$ represents the Hadamard (element-wise) product.
-Boldface lowercase and uppercase letters denote vectors and matrices.
-The trace of the matrix $\mathbf{X}$ is denoted by $\text{Tr}(\mathbf{X})$.
-$\text{diag}(\mathbf{x})$ represents the stacking of $\mathbf{x}$ on the main diagonal of a matrix.
-$\mathbf{I}_{N}$ represents an $N\times N$ identity matrix.
-The notation $\mathcal{CN}(0,\sigma^{2})$ represents circularly symmetric complex Gaussian distribution with variance $\sigma^{2}$.
-By $\mathcal{L}\setminus\{l\}$, we denote the set obtained by removing the element $l$ from the set $\mathcal{L}$.
-Given a matrix $\mathbf{M}$, $\mathbf{M}^{\dagger}$ represents its Moore-Penrose pseudoinverse.
+The contributions of this paper are outlined as follows
+  • we investigate the ability of a RIS to improve target detection under transmitter-target obstruction.
+The need for precise channel state information (CSI) for target detection and the requirement for sensing-specific streams are explored within a millimeter-wave (mmWave) channel architecture.
+  • A signal-to-clutter-and-noise ratio (SCNR)-maximizing alternating optimization (AO) algorithm is proposed, where the system assumes a line-of-sight (LoS) structure for the sensing links and optimizes the transmit power allocation and the RIS phase-shift matrix
+  • numerical simulations reveal that clutter awareness and dedicated sensing streams significantly enhance the detection probability.
 II System Model
@@
 1.
-Here a set $\mathcal{T}$ made of $T$ APs, each equipped with $M$ transmit antennas, serves $K$ single-antenna UEs in the downlink while sensing the potential presence of a target at a known position.
-The sensing is assumed to be carried out over $\tau$ symbols belonging to the same channel coherence block.
-A second set of $R$ APs, denoted by $\mathcal{R}$, collects the echos scattered from the target and performs target detection.
-The choice of geographically separated transmitting and receiving APs is motivated by the implementation challenges of full duplex operation, especially in terms of self interference [10], and the interest in this setup for cell-free massive MIMO communications.
-We assume that the direct channel between the APs in $\mathcal{T}$ and the target is blocked and, thus, the formers reach the target through a RIS equipped with $N$ reflective elements whose phase shifts are described by the vector $\boldsymbol{\theta}=[\theta_{1},\dots,\theta_{N}]^{\top}$, where $|\theta_{n}|=1$ for $n=1,\dots,N$.
+A set $\mathcal{T}$ of $T$ APs, each equipped with $M$ transmit antennas, serves $K$ single-antenna UEs in the downlink while sensing the potential presence of a target at a known position.
+Sensing is carried out over $\tau$ symbols belonging to the same channel coherence block.
+A second set $\mathcal{R}$ of $R$ APs collects the target echoes and performs target detection.
+The choice of geographically separated transmitting and receiving APs is motivated by the implementation challenges of full duplex operation, especially in terms of self-interference [8], and the interest in this setup for cell-free massive MIMO communications.
+We assume that the direct channel between the APs in $\mathcal{T}$ and the target is blocked and, thus, the formers reach the target through a RIS equipped with $N$ reflective elements, whose phase shifts are described by the vector $\boldsymbol{\theta}=[\theta_{1},\dots,\theta_{N}]^{\top}$, where $|\theta_{n}|=1$ for $n=1,\dots,N$.
 Each AP $t\in\mathcal{T}$ is connected to the RIS via the channel $\mathbf{H}_{t}\in\mathbb{C}^{N\times M}$, the reflected path connecting the RIS to the target is denoted by $\mathbf{c}\in\mathbb{C}^{N}$ and the UEs are connected to the APs in $\mathcal{T}$ via the channel $\{\mathbf{g}_{k,t}\}_{\forall k,t}\in\mathbb{C}^{M}$.
@@
 II-A Channel models
-Each of the previously mentioned channels is modeled in accordance to [11], which accurately captures the spatial sparsity that characterizes mmWave channels.
+Each of the previously mentioned channels is modelled in accordance with [9], which accurately captures the spatial sparsity that characterizes mmWave channels.
 Each channel consists of contributions from $C_{i}$ scattering clusters, i.e.,
@@
 where the overline superscript indicates the LoS paths and the other terms correspond to the NLoS parts.
-The $\beta_{0,r}^{2}$, $\beta_{0}^{2}$, and $\beta_{0,t}^{2}$ are the corresponding LoS channel gains.
+The coefficients $\beta_{0,r}^{2}$, $\beta_{0}^{2}$, and $\beta_{0,t}^{2}$ are the corresponding LoS channel gains.
 We now remove the additional subscripts and superscripts except for $n$ as the following consideration applies regardless of the additional subscripts.
@@
 The geometry of the $n$-th cluster is described by $\boldsymbol{\omega}_{n}=[\psi_{n},\phi_{n}]$, where $\psi_{n}$ is the azimuth angle and $\phi_{n}$ is the elevation angle.
-Each AP is equipped with a uniform planar square array with a half-wavelength vertical and horizontal inter-element spacing such that the entries of $M$-dimensional steering vector are defined as
-$[\mathbf{a}_{M}(\boldsymbol{\omega}_{n})]_{m_{h}m_{v}}=e^{-j\pi(m_{h}\sin(\psi_{n})\cos(\phi_{n})+m_{v}\sin(\phi_{n}))}$ (5)
-for $m_{h},m_{v}=0,\dots,\sqrt{M}-1$.
-We assume that perfect CSI for $\{\mathbf{g}_{k,t}\}_{\forall k,t}$ is available at the APs, as this is easily obtainable through well-known estimation techniques.
-As for the sensing channels, the transmit APs only know the position of the RIS and the target and, thus, the corresponding LoS channels.
+Each AP is equipped with a uniform planar square array with a half-wavelength vertical and horizontal inter-element spacing such that the entries of $M$-dimensional steering vector are defined as $[\mathbf{a}_{M}(\boldsymbol{\omega}_{n})]_{m_{h}m_{v}}=e^{-j\pi(m_{h}\sin(\psi_{n})\cos(\phi_{n})+m_{v}\sin(\phi_{n}))}$,for $m_{h},m_{v}=0,\dots,\sqrt{M}-1$.
+We assume that perfect CSI regarding $\{\mathbf{g}_{k,t}\}_{\forall k,t}$ is available at the APs, as this is easily obtainable through standard estimation techniques.
+As for the sensing channels, the transmit APs know the position of the RIS and the target and, thus, the corresponding LoS channels.
 II-B Communication observation model
@@
 To this end, we define $\mathbf{g}_{k}=[\mathbf{g}_{k,1}^{\top},\dots,\mathbf{g}_{k,T}^{\top}]^{\top}\in\mathbb{C}^{MT}$ and $\mathbf{f}_{k}=[\mathbf{f}_{k,1}^{\top},\dots,\mathbf{f}_{k,T}^{\top}]^{\top}\in\mathbb{C}^{MT}$.
-The communication precoding vectors for each UE are chosen using RZF as
-$\mathbf{f}_{k}=\frac{\left(\sum_{i\in\mathcal{K}}\mathbf{g}_{i}\mathbf{g}_{i}^{\mathsf{H}}+\lambda\mathbf{I}_{MT}\right)^{-1}\mathbf{g}_{k}}{\left\lVert\left(\sum_{i\in\mathcal{K}}\mathbf{g}_{i}\mathbf{g}_{i}^{\mathsf{H}}+\lambda\mathbf{I}_{MT}\right)^{-1}\mathbf{g}_{k}\right\rVert},\leavevmode\nobreak\ k=1,\dots,K,$ (7)
-where $\lambda$ is a regularization parameter and this parameter allows the system to implement a trade-off between interference and noise suppression.
-For the sensing stream, let us define the matrix $\mathbf{G}=[\mathbf{g}_{1},\dots,\mathbf{g}_{K}]\in\mathbb{C}^{MT\times K}$ with all the UE channels.
-Then the sensing stream precoder is defined as
-$\mathbf{f}_{L,t}=\frac{\left(\mathbf{I}_{MT}-\mathbf{G}\left(\mathbf{G}^{\mathsf{H}}\mathbf{G}\right)^{\dagger}\mathbf{G}^{\mathsf{H}}\right)\overline{\mathbf{h}}}{\left\lVert\left(\mathbf{I}_{MT}-\mathbf{G}\left(\mathbf{G}^{\mathsf{H}}\mathbf{G}\right)^{\dagger}\mathbf{G}^{\mathsf{H}}\right)\overline{\mathbf{h}}\right\rVert},$ (8)
+The communication precoding vectors for each UE are chosen using regularized zero forcing (RZF) as $\mathbf{f}_{k}=\overline{\mathbf{f}}_{k}/\left\lVert\overline{\mathbf{f}}_{k}\right\rVert$, where
+$\overline{\mathbf{f}}_{k}=\left(\sum_{i\in\mathcal{K}}\mathbf{g}_{i}\mathbf{g}_{i}^{\mathsf{H}}+\lambda\mathbf{I}_{MT}\right)^{-1}\mathbf{g}_{k}$ (6)
+where $\lambda$ is a regularization parameter.
+Let us define the matrix $\mathbf{G}=[\mathbf{g}_{1},\dots,\mathbf{g}_{K}]\in\mathbb{C}^{MT\times K}$ with all the UE channels.
+Then the sensing stream precoder is defined as $\mathbf{f}_{L,t}=\overline{\mathbf{f}}_{L,t}/\left\lVert\overline{\mathbf{f}}_{L,t}\right\rVert$
+$\overline{\mathbf{f}}_{L,t}=\left(\mathbf{I}_{MT}-\mathbf{G}\left(\mathbf{G}^{\mathsf{H}}\mathbf{G}\right)^{\dagger}\mathbf{G}^{\mathsf{H}}\right)\overline{\mathbf{h}}$ (7)
 where $\overline{\mathbf{h}}=[\overline{\mathbf{h}}_{1}^{\top},\dots,\overline{\mathbf{h}}_{T}^{\top}]^{\top}$ and $\overline{\mathbf{h}}_{t}=\overline{\mathbf{H}}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\overline{\mathbf{c}}$ with $\boldsymbol{\Theta}=\text{diag}(\boldsymbol{\theta})$.
-We have projected $\overline{\mathbf{h}}$ onto the null space of the communication UEs’ channels to null the interference caused to the UEs.
-The communication performance is represented by the UEs’ spectral efficiency (SE), defined as ${\rm{SE}}_{k}=\log_{2}(1+{\rm{SINR}}_{k})$: under the assumption of perfect CSI on the UEs channels, the SINR is defined as
+We have projected $\overline{\mathbf{h}}$ onto the null space of the communication UEs’ channels to nullify the interference caused to the UEs.
+The communication performance is represented by the UEs’ spectral efficiency (SE), defined as ${\rm{SE}}_{k}=\log_{2}(1+{\rm{SINR}}_{k})$: under the assumption of perfect CSI on the UEs channels, the signal-to-interference-plus-noise ratio (SINR) is defined as
 ${\rm{SINR}}_{k}=\frac{|{\mathbf{g}}_{k}^{\mathsf{H}}\mathbf{f}_{k}\sqrt{\rho_{k}}|^{2}}{\sum_{k^{\prime}\in\mathcal{K}/k}|{\mathbf{g}}_{k}^{\mathsf{H}}\mathbf{f}_{k^{\prime}}\sqrt{\rho_{k^{\prime}}}|^{2}+|{\mathbf{g}}_{k}^{\mathsf{H}}\mathbf{f}_{L}\sqrt{\rho_{L}}|^{2}+\sigma_{k}^{2}}.$ (8)
@@
 The two-way sensing channel between transmit AP $t$ and receive AP $r$ is defined as $\mathbf{E}_{t,r}=\mathbf{b}_{r}\mathbf{h}_{t}^{\mathsf{H}}\triangleq\mathbf{b}_{r}\left(\mathbf{H}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\mathbf{c}\right)^{\mathsf{H}}.$ The sensing observation of AP $r$ at timeslot $\iota$ is
-$\mathbf{y}_{r}[\iota]=\sum_{t\in\mathcal{T}}\xi_{t,r}\mathbf{E}_{t,r}\mathbf{s}_{t}[\iota]+\mathbf{z}_{r}[\iota]+\mathbf{n}_{r}[\iota],$ (10)
+$\mathbf{y}_{r}[\iota]=\sum_{t\in\mathcal{T}}\xi_{t,r}\mathbf{E}_{t,r}\mathbf{s}_{t}[\iota]+\mathbf{z}_{r}[\iota]+\mathbf{n}_{r}[\iota],\vspace{-2mm}$ (9)
 where $\xi_{t,r}$ is the radar cross section (RCS) associated with the sensing path between transmit AP $t$ and receive AP $r$.
 We adopt the Swerling-I model, meaning that the RCS assumes only one value during the transmission and $\xi_{t,r}\sim\mathcal{CN}(0,\delta_{t,r}^{2})$.
-We assume that $\{\delta_{t,r}^{2}\}_{\forall t,r}$ are known and that RCSs belonging to different transmit-receiving pairs are statistically uncorrelated, that is
-$\mathbb{E}\left\{\xi_{a,r}\xi_{a^{\prime},r^{\prime}}^{*}\right\}=\begin{cases}\delta_{a,r}^{2},&{\rm{if}}\leavevmode\nobreak\ a=a^{\prime},\leavevmode\nobreak\ r=r^{\prime},\\ 0,&\rm{otherwise}.\end{cases}$ (11)
+We assume that $\{\delta_{t,r}^{2}\}_{\forall t,r}$ are known and that RCSs belonging to different transmit-receiving pairs are statistically uncorrelated, that is $\mathbb{E}\left\{\xi_{a,r}\xi_{a^{\prime},r^{\prime}}^{*}\right\}\neq 0\Leftrightarrow a=a^{\prime},\leavevmode\nobreak\ r=r^{\prime}$.
 In (9), $\mathbf{n}_{r}[\iota]$ is the receiver noise of AP $r$, assumed to have independent $\mathcal{CN}(0,\sigma^{2})$ entries.
 The vector $\mathbf{z}_{r}[\iota]$ represents the clutter as observed from AP $r$.
-The clutter vector is modeled as $\mathbf{z}_{r}[\iota]\sim\mathcal{CN}(\boldsymbol{0},\delta_{\mathbf{z}}^{2}\mathbf{R}_{r})$, where $\delta_{\mathbf{z}}^{2}$ is the power of the clutter cross section, which also includes the path loss effects, thus defining the CNR as $\textrm{CNR}=\delta_{\mathbf{z}}^{2}/\sigma^{2}$.
-The normalized clutter covariance matrix with trace being equal to $M$ is denoted by $\mathbf{R}_{r}$.
+The clutter vector is modelled as $\mathbf{z}_{r}[\iota]\sim\mathcal{CN}(\boldsymbol{0},\delta_{\mathbf{z}}^{2}\mathbf{R}_{r})$, where $\delta_{\mathbf{z}}^{2}$ is the power of the clutter cross-section, which also includes the path loss effects, thus defining the clutter-to-noise ratio (CNR) as $\textrm{CNR}=\delta_{\mathbf{z}}^{2}/\sigma^{2}$.
+The normalized clutter covariance matrix, with a trace equal to $M$, is denoted by $\mathbf{R}_{r}$.
 This matrix is unknown, but its eigenspace is assumed to be known later in the detector design.
@@
 $\overline{\textrm{SCNR}}=\frac{\mathbb{\sum}_{\iota=1}^{\tau}\mathbb{\sum}_{r\in\mathcal{R}}\sum_{a\in\mathcal{T}}\delta_{a,r}^{2}\left\lVert\overline{\mathbf{E}}_{a,r}\mathbf{s}_{a}[\iota]\right\rVert^{2}}{R\tau M\left(\sigma^{2}+\delta_{\mathbf{z}}^{2}\right)},$ (10)
-where $\overline{\mathbf{E}}_{a,r}=\overline{\mathbf{b}}_{r}\left(\overline{\mathbf{H}}_{a}^{\mathsf{H}}\boldsymbol{\Theta}\overline{\mathbf{c}}\right)^{\mathsf{H}}$
+where $\overline{\mathbf{E}}_{a,r}=\overline{\mathbf{b}}_{r}\left(\overline{\mathbf{H}}_{a}^{\mathsf{H}}\boldsymbol{\Theta}\overline{\mathbf{c}}\right)^{\mathsf{H}}$.
 III Sensing performance optimization
@@
 III-A Power allocation optimization
-In this subsection, we introduce the first subproblem of the AO algorithm, where we optimize the power allocation for a fixed RIS configuration. We first reformulate (12) as
-$\overline{\textrm{SCNR}}=\boldsymbol{\rho}^{\top}\Re(\mathbf{A})\boldsymbol{\rho},$ (14)
-where
+In this subsection, we introduce the first subproblem of the AO algorithm, where we optimize the power allocation for a fixed RIS configuration.
+We first reformulate (10) as $\overline{\textrm{SCNR}}=\boldsymbol{\rho}^{\top}\Re(\mathbf{A})\boldsymbol{\rho}$ where
 $\mathbf{A}=\frac{\sum_{\iota=1}^{\tau}\mathbb{\sum}_{r\in\mathcal{R}}\sum_{a\in\mathcal{T}}\delta_{a,r}^{2}\mathbf{X}^{\mathsf{H}}[\iota]\mathbf{F}_{a}^{\mathsf{H}}\overline{\mathbf{E}}_{a,r}^{\mathsf{H}}\overline{\mathbf{E}}_{a,r}\mathbf{F}_{a}\mathbf{X}[\iota]}{R\tau M\left(\sigma^{2}+\delta_{\mathbf{z}}^{2}\right)}.$ (12)
-The set of constraints (13b) can be reformulated as second order cone (SOC) constraints in terms of $\boldsymbol{\rho}$ as
-$\left\lVert\begin{bmatrix}|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{1}|\sqrt{\rho_{1}}&\dots&\underbrace{0}_{k}&\dots&|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{L}|\sqrt{\rho_{L}}&\sigma_{k}\end{bmatrix}\right\rVert$
-$\leq\frac{|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{k}|\sqrt{\rho_{k}}}{\sqrt{\gamma_{k}}}.$ (16)
-The power constraint in (13c) can also be rewritten in SOC form as
-$\left\lVert\text{diag}\left(\left\lVert\mathbf{f}_{1,t}\right\rVert,\dots,\left\lVert\mathbf{f}_{L,t}\right\rVert\right)\boldsymbol{\rho}\right\rVert\leq\sqrt{P_{t}}.$ (17)
-The first subproblem of the AO SCNR-maximizing algorithm is defined as
+The constraints (11b) and (11c) can be reformulated as second order cone (SOC) constraints in terms of $\boldsymbol{\rho}$. The first subproblem of the AO SCNR-maximizing algorithm is defined as
 $\underset{\boldsymbol{\rho}}{\mathrm{maximize}}\,\,$ $\leavevmode\nobreak\ \boldsymbol{\rho}^{\top}\Re(\mathbf{A})\boldsymbol{\rho}$ (13a)
-[…]
+subject to $\Big{[}|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{1}|\sqrt{\rho_{1}}\quad\dots\quad\underbrace{0}_{k}\quad\dots\quad|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{L}|\sqrt{\rho_{L}}$
+$\sigma_{k}\Big{]}\leq(|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{k}|\sqrt{\rho_{k}})/{\sqrt{\gamma_{k}}},\leavevmode\nobreak\ \forall k\in\mathcal{K}$ (13b)
+$\left\lVert\widetilde{\mathbf{F}}_{t}\boldsymbol{\rho}\right\rVert\leq\sqrt{P_{t}},\leavevmode\nobreak\ \forall t\in\mathcal{T}$ (13c)
+where $\widetilde{\mathbf{F}}_{t}=\text{diag}\left(\left\lVert\mathbf{f}_{1,t}\right\rVert,\dots,\left\lVert\mathbf{f}_{L,t}\right\rVert\right)$.
 This problem can be solved using the convex-concave procedure outlined in [3].
 III-B RIS phase-shift optimization
-We now move on to the second part of the AO, where we optimize $\boldsymbol{\theta}$ when the power variables are fixed. The SCNR can be reformulated as
-$\overline{\textrm{SCNR}}=\boldsymbol{\theta}^{\mathsf{H}}\mathbf{Q}\boldsymbol{\theta},$ (19)
-where
+We now move on to the second part of the AO, where we optimize $\boldsymbol{\theta}$ when the power variables are fixed.
+The SCNR can be reformulated as $\overline{\textrm{SCNR}}=\boldsymbol{\theta}^{\mathsf{H}}\mathbf{Q}\boldsymbol{\theta}$, where
 $\mathbf{Q}=$ $\frac{1}{R\tau M\left(\sigma^{2}+\delta_{\mathbf{z}}^{2}\right)}\sum_{\iota=1}^{\tau}\sum_{r\in\mathcal{R}}\sum_{a\in\mathcal{T}}\Big{(}\delta_{a,r}^{2}\overline{\mathbf{c}}\overline{\mathbf{b}}_{r}^{\mathsf{H}}\overline{\mathbf{b}}_{r}\overline{\mathbf{c}}^{\mathsf{H}}\Big{)}^{\top}$
 $\quad\quad\odot\left(\overline{\mathbf{H}}_{a}\mathbf{s}_{a}[\iota]\mathbf{s}_{a}^{\mathsf{H}}[\iota]\overline{\mathbf{H}}_{a}^{\mathsf{H}}\right).$ (14)
-We have used $\text{Tr}(\mathbf{A}\mathbf{X}\mathbf{B}\mathbf{X}^{\mathsf{H}})=\text{vec}^{\mathsf{H}}(\mathbf{X})\left(\mathbf{B}^{\top}\otimes\mathbf{A}\right)\text{vec}(\mathbf{X})$ and the following lemma.
-Lemma 1.
-Let $\mathbf{x}\in\mathbb{C}^{N}$ and $\mathbf{X}=\text{diag}(\mathbf{x})$. For any $\mathbf{A}\in\mathbb{C}^{N\times N}$ and $\mathbf{B}\in\mathbb{C}^{N\times N}$, it holds that
-$\text{vec}^{\mathsf{H}}(\mathbf{X})(\mathbf{A}\otimes\mathbf{B})\text{vec}(\mathbf{X})=\mathbf{x}^{\mathsf{H}}(\mathbf{A}\odot\mathbf{B})\mathbf{x}.$ (21)
-Proof: Let $\text{vec}(\mathbf{X})=\sum_{n=1}^{N}x_{n}\mathbf{e}_{n}\otimes\mathbf{e}_{n}$, where $\mathbf{e}_{n}$ is the $n$-th column of $\mathbf{I}_{N}$.
-Then, the right-hand side of (21) can be written as $\left(\sum_{n=1}^{N}x_{n}\mathbf{e}_{n}\otimes\mathbf{e}_{n}\right)^{\mathsf{H}}(\mathbf{A}\otimes\mathbf{B})\left(\sum_{m=1}^{N}x_{m}\mathbf{e}_{m}\otimes\mathbf{e}_{m}\right)\overset{(a)}{=}\sum_{n=1}^{N}\sum_{m=1}^{N}x_{n}x_{m}^{\mathsf{H}}(\mathbf{e}_{n}^{\mathsf{H}}\mathbf{A}\mathbf{e}_{m})(\mathbf{e}_{n}^{\mathsf{H}}\mathbf{B}\mathbf{e}_{m})=\sum_{n=1}^{N}\sum_{m=1}^{N}x_{n}x_{m}^{\mathsf{H}}[\mathbf{A}]_{n,m}[\mathbf{B}]_{n,m}=\mathbf{x}^{\mathsf{H}}(\mathbf{A}\odot\mathbf{B})\mathbf{x}$, where in $(a)$ we used mixed-product property of the Kronecker product.
-The RIS phase-shift optimization problem under consideration is still non-convex.
-We are trying to maximize a quadratic form of $\boldsymbol{\theta}$.
-We then employ the minorization-maximization (MM) algorithm to maximize a convex lower bound on (19) [12].
-More specifically, a convex local lower bound can be found by applying the successive convex approximation [13] around the local point $\boldsymbol{\theta}^{(s)}$ at iteration $s$.
-The objective function can be lower bounded by
-$\boldsymbol{\theta}^{\mathsf{H}}\mathbf{Q}\boldsymbol{\theta}\geq 2\Re\left(\boldsymbol{\theta}^{(s)\mathsf{H}}\mathbf{Q}^{(+)}\boldsymbol{\theta}\right)+\boldsymbol{\theta}^{\mathsf{H}}\mathbf{Q}^{(-)}\boldsymbol{\theta}-\boldsymbol{\theta}^{(s)\mathsf{H}}\mathbf{Q}^{(+)}\boldsymbol{\theta}^{(s)},$ (22)
-where $\mathbf{Q}^{(+)}$ and $\mathbf{Q}^{(-)}$ represent the positive and negative semidefinite parts of $\mathbf{Q}$. We can now define the second subproblem of the AO optimization as
+We have used $\text{Tr}(\mathbf{A}\mathbf{X}\mathbf{B}\mathbf{X}^{\mathsf{H}})=\text{vec}^{\mathsf{H}}(\mathbf{X})\left(\mathbf{B}^{\top}\otimes\mathbf{A}\right)\text{vec}(\mathbf{X})$ and $\text{vec}^{\mathsf{H}}(\text{diag}(\mathbf{x}))(\mathbf{A}\otimes\mathbf{B})\text{vec}(\text{diag}(\mathbf{x}))=\mathbf{x}^{\mathsf{H}}(\mathbf{A}\odot\mathbf{B})\mathbf{x}$ from [10].
+The RIS phase-shift optimization problem under consideration is still non-convex: We then employ the minorization-maximization (MM) algorithm to maximize a convex lower bound on the objective function [11] around the local point $\boldsymbol{\theta}^{(s)}$.
+We can now define the second subproblem of the AO optimization as
 $\underset{\boldsymbol{\theta},\alpha\geq 0}{\textrm{maximize}}$ $\leavevmode\nobreak\ \alpha$ (15a)
@@
 $|\theta_{n}|\leq 1,\quad n=1,\dots,N,$ (15c)
-where $\alpha$ is an auxiliary variable.
+where $\alpha$ is an auxiliary variable and $\mathbf{Q}^{(+)}$ and $\mathbf{Q}^{(-)}$ represent the positive and negative semidefinite parts of $\mathbf{Q}$.
 Note that we have relaxed the unit modulus constraint, as it is observed that the solution satisfies this constraint with equality anyway.
@@
 IV GLRT Detector
-One of the main metrics used to assess the sensing performance is the probability of correct detection, implemented through a GLRT by extending the detector in [9] to a cell-free massive MIMO system based on the optimization procedure developed in the previous section.
+One of the main metrics used to assess the sensing performance is the probability of correct detection, implemented through a generalized likelihood ratio test (GLRT) by extending the detector in [7] to a cell-free massive MIMO system based on the optimization procedure developed in the previous section.
 We operate under the assumption that the transmit signals are known by the receiving APs, thanks to the cell-free architecture.
 We define the vector containing the combined sensing observation of all receiving APs during timeslot $\iota$ as
-$\mathbf{y}[\iota]=[\mathbf{y}_{1}^{\top}[\iota],\dots,\mathbf{y}_{R}^{\top}[\iota]]^{\top}=\mathbf{V}[\iota]\boldsymbol{\xi}+\mathbf{U}\mathbf{w}[\iota]+\mathbf{n}[\iota],$ (24)
-$\mathbf{V}[\iota]=\textrm{blkdiag}(\mathbf{V}_{1}[\iota],\dots,\mathbf{V}_{R}[\iota])\in\mathbb{C}^{MR\times TR},$ (25)
-$\mathbf{V}_{r}[\iota]=\left[\mathbf{E}_{1,r}\mathbf{s}_{1}[\iota],\dots,\mathbf{E}_{T,r}\mathbf{s}_{T}[\iota]\right]\in\mathbb{C}^{M\times T},$ (26)
-$\mathbf{n}[\iota]=[\mathbf{n}_{1}^{\top}[\iota],\dots,\mathbf{n}_{R}^{\top}[\iota]]^{\top},\leavevmode\nobreak\ \boldsymbol{\xi}=[\xi_{1,1},\dots,\xi_{T,R}]^{\top},$ (27)
-$\mathbf{U}=\textrm{blkdiag}\left(\mathbf{U}_{1},\dots,\mathbf{U}_{R}\right),\mathbf{w}[\iota]=[\mathbf{w}_{1}^{\top}[\iota],\dots,\mathbf{w}_{R}^{\top}[\iota]]^{\top}.$ (28)
+$\mathbf{y}[\iota]=[\mathbf{y}_{1}^{\top}[\iota],\dots,\mathbf{y}_{R}^{\top}[\iota]]^{\top}=\mathbf{V}[\iota]\boldsymbol{\xi}+\mathbf{U}\mathbf{w}[\iota]+\mathbf{n}[\iota],$
+$\mathbf{V}[\iota]=\textrm{blkdiag}(\mathbf{V}_{1}[\iota],\dots,\mathbf{V}_{R}[\iota])\in\mathbb{C}^{MR\times TR},$
+$\mathbf{V}_{r}[\iota]=\left[\mathbf{E}_{1,r}\mathbf{s}_{1}[\iota],\dots,\mathbf{E}_{T,r}\mathbf{s}_{T}[\iota]\right]\in\mathbb{C}^{M\times T},$
+$\mathbf{n}[\iota]=[\mathbf{n}_{1}^{\top}[\iota],\dots,\mathbf{n}_{R}^{\top}[\iota]]^{\top},\leavevmode\nobreak\ \boldsymbol{\xi}=[\xi_{1,1},\dots,\xi_{T,R}]^{\top},$
+$\mathbf{U}=\textrm{blkdiag}\left(\mathbf{U}_{1},\dots,\mathbf{U}_{R}\right),\mathbf{w}[\iota]=[\mathbf{w}_{1}^{\top}[\iota],\dots,\mathbf{w}_{R}^{\top}[\iota]]^{\top}.$
 The clutter observed by AP $r$ is assumed to exist in a low-rank subspace [12] spanned by the columns of the semi-unitary matrix $\mathbf{U}_{r}\in\mathbb{C}^{M\times i}$, with $i$ being the subspace dimension, obtained from the eigenspace of $\delta_{\mathbf{z}}^{2}\mathbf{R}_{r}$.
@@
 where $\overline{\mathbf{V}}$ is computed as $\mathbf{V}$ by using $\overline{\mathbf{E}}_{t,r}$.
-The adoption of a GLRT method is motivated by the unknown distribution of $\mathbf{w}[\iota]$, however we know that $\mathbf{y}_{\mathcal{H}_{0}}[\iota]-\mathbf{U}\mathbf{w}[\iota]\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{MR})$ and $\mathbf{y}_{\mathcal{H}_{1}}[\iota]-\mathbf{U}\mathbf{w}[\iota]\sim\mathcal{CN}(\mathbf{0},\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR})$, where $\mathbf{R}[\iota]=\mathbb{E}\left\{\overline{\mathbf{V}}[\iota]\boldsymbol{\xi}\boldsymbol{\xi}^{\mathsf{H}}\overline{\mathbf{V}}^{\mathsf{H}}[\iota]\right\}=\overline{\mathbf{V}}[\iota]\text{diag}(\delta^{2}_{1,1},\dots,\delta^{2}_{T,R})\overline{\mathbf{V}}^{\mathsf{H}}[\iota]$.
-The GLRT test is then defined as
-$\frac{\prod_{\iota=1}^{\tau}\max_{\mathbf{w}[\iota]}f_{\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}_{1}}\left(\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}_{1}\right)}{\prod_{\iota=1}^{\tau}\max_{\mathbf{w}[\iota]}f_{\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}_{0}}\left(\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}_{0}\right)}\lessgtr^{\mathcal{H}_{0}}_{\mathcal{H}_{1}}\lambda_{d},$ (30)
-where $f_{\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}}$ denote the probability density function. Following the derivation in [9], the GLRT test can be rewritten as
-$\sum_{\iota=1}^{\tau}\mathbf{y}^{\mathsf{H}}[\iota]\mathbf{T}[\iota]\mathbf{y}[\iota]\lessgtr^{\mathcal{H}_{0}}_{\mathcal{H}_{1}}\underbrace{\ln(\lambda_{d})-\ln\left(\varsigma\right)}_{\widetilde{\lambda}_{d}}$ (31)
-$\mathbf{T}[\iota]=\boldsymbol{\Xi}[\iota]\mathbf{U}\Big{(}\mathbf{U}^{\mathsf{H}}\boldsymbol{\Xi}[\iota]\mathbf{U}\Big{)}^{-1}\mathbf{U}^{\mathsf{H}}\boldsymbol{\Xi}[\iota]$
-$\quad\quad+\sigma^{-2}(\mathbf{I}_{MR}-\mathbf{U}\mathbf{U}^{\mathsf{H}})-\boldsymbol{\Xi}[\iota]$ (32)
-$\boldsymbol{\Xi}[\iota]=(\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR})^{-1}$ (33)
-where $\varsigma=\det\left(\sigma^{2}\mathbf{I}_{MR}\right)^{\tau}/{\prod_{\iota=1}^{\tau}\det\left(\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR}\right)}$.
-$\widetilde{\lambda}_{d}$ is a revised threshold whose value should be chosen according to the desired false alarm probability, and it can be computed through Monte Carlo trials.
+The adoption of a GLRT method is motivated by the unknown distribution of $\mathbf{w}[\iota]$, however, we know that $\mathbf{y}_{\mathcal{H}_{0}}[\iota]-\mathbf{U}\mathbf{w}[\iota]\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{MR})$ and $\mathbf{y}_{\mathcal{H}_{1}}[\iota]-\mathbf{U}\mathbf{w}[\iota]\sim\mathcal{CN}(\mathbf{0},\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR})$, where $\mathbf{R}[\iota]=\mathbb{E}\left\{\overline{\mathbf{V}}[\iota]\boldsymbol{\xi}\boldsymbol{\xi}^{\mathsf{H}}\overline{\mathbf{V}}^{\mathsf{H}}[\iota]\right\}=\overline{\mathbf{V}}[\iota]\text{diag}(\delta^{2}_{1,1},\dots,\delta^{2}_{T,R})\overline{\mathbf{V}}^{\mathsf{H}}[\iota]$.
+Following the derivation in [7], the GLRT test can be rewritten as
+$\sum_{\iota=1}^{\tau}\mathbf{y}^{\mathsf{H}}[\iota]\Big{(}\boldsymbol{\Xi}[\iota]\mathbf{U}\Big{(}\mathbf{U}^{\mathsf{H}}\boldsymbol{\Xi}[\iota]\mathbf{U}\Big{)}^{-1}\mathbf{U}^{\mathsf{H}}\boldsymbol{\Xi}[\iota]+$
+$\sigma^{-2}(\mathbf{I}_{MR}-\mathbf{U}\mathbf{U}^{\mathsf{H}})-\boldsymbol{\Xi}[\iota]\Big{)}\mathbf{y}[\iota]\lessgtr^{\mathcal{H}_{0}}_{\mathcal{H}_{1}}\widetilde{\lambda}_{d}$ (17)
+where $\boldsymbol{\Xi}[\iota]=(\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR})^{-1}$.
+$\widetilde{\lambda}_{d}=\ln(\lambda_{d})-\ln\left(\det\left(\sigma^{2}\mathbf{I}_{MR}\right)^{\tau}/{\prod_{\iota=1}^{\tau}\det\left(\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR}\right)}\right)$ is a revised threshold whose value should be chosen according to the desired false alarm probability.
 5050505010010010010015015015015000101010102020202030303030x (m)y (m)$\mathcal{T}$$\mathcal{R}$
@@
 V Numerical Results
-In this section, the numerical simulation results are provided.
-A two-dimensional view of the simulated scenario is provided in Fig.
-2.
-The locations of the transmit APs in $\mathcal{T}$, the receiving ones in $\mathcal{R}$, and the UEs are randomly generated within the respective bounding boxes, and their $z$ coordinates are equal to $10$ m, $10$ m, and $1$ m, respectively.
-The RIS and target position are assumed to be fixed and their $z$ coordinates are equal to $15$ m and $3$ m.
-The power budget for each transmit AP is $2$ dBW, with a bandwidth of $B=1$ MHz, the noise power is equal to $-204+10\log_{10}(B)$ dBW.
-The channel’s path loss follows the 3GPP urban microcell model defined in [15].
-The clutter covariance matrix $\mathbf{R}_{r}$ is modeled according to the local scattering model in [16] with six clusters centered around the $\boldsymbol{\omega}_{0,r}^{\textrm{RX}}\pm[20^{\circ}$, $10^{\circ}]$.
-We have $T=5$ transmit APs and $R=2$ receiving APs, each equipped with $M=36$ antennas and serving $K=5$ UEs.
-We want to guarantee an SINR of $\gamma_{k}=3$ dB to all UEs.
-Unless otherwise specified, the RIS is equipped with $N=64$ elements, $C_{1}=C_{2}=C_{3}=C_{4}=2$, and the angles are uniformly distributed within an angle spread of $\sigma_{\psi_{n}}=10^{\circ}$ in azimuth and $\sigma_{\phi_{n}}=10^{\circ}$ in elevation around the LoS angles between transmitter and receiver, denoted by $\psi_{0}\leavevmode\nobreak\ (\phi_{0})$.
-The results shown hereafter are averaged over $10$ random realizations of the AP positions and UE positions.
-For each of these realizations, the probability of detection is computed over $\tau=5$ timeslots and $1000\tau$ noise and clutter realizations.
-The false alarm probability is set to $10^{-4}$.
 […]
-Figure 3: Probability of detection with respect to the RCS power with a CNR of $20$ dB.
 […]
-Figure 4: Probability of detection with respect to the RCS power with a CNR of $40$ dB.
+Figure 3: Probability of detection vs RCS power with CNR = $20$ dB.
+[…]
+Figure 4: Probability of detectionvs the RCS power with a CNR=$40$ dB.
 V-A Clutter awareness and cluster number’s impact
-We first analyze the impact of the clutter awareness of the receiving APs on the detection probability and how much the LoS assumed structure of the channels affects the system’s detection performances when there are many clusters concentrated within a small angular region.
 Fig.
@@
 the RCS power, where we assumed that $\delta_{1,1}^{2}=\dots=\delta_{T,R}^{2}=\delta^{2}$ for a CNR of $20$ dB.
-We can see that a clutter-aware system outperforms a clutter-unaware one: The clutter unaware dectector is obtained from (29) by removing the term $\mathbf{U}\mathbf{w}[\iota]$, thus obtaining $\mathbf{T}[\iota]=\sigma^{-2}\mathbf{I}_{MR}-\boldsymbol{\Xi}[\iota]$.
-On the other hand, when $C_{2}=C_{3}=C_{4}=20$, here denoted by $C_{234}$, the LoS assumption does not hold as much, thus generating a performance decrease.
-Our optimized model greatly outperforms a benchmark approach, consisting of a random RIS phase-shift matrix and an equal power allocation between the beams.
+We can see that a clutter-aware system outperforms a clutter-unaware one: The clutter-unaware detector is defined as $\mathbf{T}[\iota]=\sigma^{-2}\mathbf{I}_{MR}-\boldsymbol{\Xi}[\iota]$.
+When $C_{2}=C_{3}=C_{4}=20$, denoted by $C_{234}$, the LoS assumption is no longer realistic, thus generating a performance decrease.
+Our optimized model greatly outperforms the benchmark approach, consisting of a random RIS phase-shift matrix and an equal power allocation between the beams.
 Fig.
@@
 […]
-Figure 5: Impact of a dedicated sensing stream onto the probability of detection with a CNR of $20$ dB.
+Figure 5: Impact of a dedicated sensing stream onto the probability of detection with a CNR=$20$ dB.
 V-B Do we need dedicated sensing symbols?
-Given the sparse nature of the mmWave channels, one might question whether or not a dedicated sensing stream is necessary and to what extent the system benefits from it.
-To gain more insight into this, we have analyzed a system without sensing-dedicated precoding.
+We have analyzed a system without sensing-dedicated precoding.
 Fig.
-5 compares the detection performance of systems with and without dedicated sensing precoding, both in the presence and in the absence of clutter awareness.
+5 compares the detection performance of systems with and without dedicated sensing precoding, both in the presence and absence of clutter awareness.
 We can immediately see that removing the sensing streams implies a performance degradation.
@@
 We have investigated a cell-free massive MIMO ISAC network jointly serving multiple communication UEs while detecting a potential target through an RIS in a multi-static fashion.
-The system wants to maximize its sensing performance while preserving a minimum SINR for the communication UEs.
-To this end, we optimize the power allocation of a precoding codebook jointly with the RIS phase-shift matrix and we derive the sufficient statistics for the target detector.
-Numerical results show that clutter awareness at the receiving APs plays a crucial role in the system’s detection performance and that including a dedicated sensing stream benefits the system.
-We see that the mmWave channel structure allows the system to assume LoS channels and still achieve good performances when the channel is composed of a high number of clusters concentrated in a small angular range.
+To this end, the transmit power allocation and the RIS phase-shift matrix are optimized to increase the target detection probability.
+Numerical results show that clutter awareness plays a crucial role in the detection performance and that including a dedicated sensing stream benefits the system.
+The mmWave channel structure allows the system to assume LoS channels and still achieve good performances when the channel is composed of a high number of clusters concentrated in a small angular range.